A 21.0 kg body is moving in the direction of the pQ5
Answer #1

a)M = 21 Kg

V = 100 i m/s

Initial momentum = (21*100) i Kgm/s

m1 = 7 Kg, v1 = 180 j m/s

m2 = 4 Kg, v2 = 250 i m/s

m3 = M-m1-m2 = 21-11 = 10 Kg

Momentum is conserved

=>(21*100)i = -(4*250)i+(7*180)j+(8*V')

=> 2100i = -1000i + 1260j + 8V' => 3100i +1260j =8V'

=>V' = 387.5i - 157.5 j

b)Magnitude = sqrt((157.5^2)+(387.5^2)) = 418.285 m/s

Direction: 22.12 degrees clockwise to the +ve x-axis

c)Change in mechanical energy = (0.5*21*(100^2))-(0.5*7*(180^2))-(0.5*4*(250^2))-(0.5*8*(418.285^2)) = 803249.36 J